# Furstenberg set problem and exceptional set estimate in prime fields: dimension two implies higher dimensions

### v1

@@ 2.2. Useful lemmas
To show this, we just need to prove for a fixed $a\in(0,n]$, the ranges of $s$ determined by four cases form a partition of $(0,\infty)$. Type 1 corresponds to $s\in(\min\{a,k\},\infty)$. Type 4 corresponds to $s\in(0,a-(n-k)]$. Note that $a-(n-k)\leq\min\{a,k\}$, so we just need to show the ranges of $s$ determined by Type 2 and Type 3 from a partition of $(a-(n-k),\min\{a,k\}]$. Since the ranges of $s$ determined by Type2 and Type 3 are disjoint, we just need to show
[…]
We first show “"absent"$\subset$"⊂ ". Suppose $s$ satisfies the condition in Type 2. We want to show $s>a-(n-k)$, which is equivalent to $m<n+l-k+\gamma-\beta$. This is true since $m\leq n+l-k$ and $\gamma>\beta$. Suppose $s$ satisfies condition in Type 3. Again, we want to show $m<n+l-k+\gamma-\beta$. This is true since $m\leq n+l-k-1$ and $\gamma-\beta>-\beta\geq-1$.
Next, we show $\supset$”. Suppose $s=l+\gamma\in(a-(n-k),\min\{a,k\}]$. $s>a-(n-k)$ implies
@@
Lemma 23.
@@
Let $a=m+\beta,s=l+\gamma$ be the canonical expression. We just check four types as in Definition 14.
@@
Note that the integer part $(m,l)$ is reduced to $(n-1,l)$, so we can keep doing the iteration.
@@
Lemma 24.
Suppose $a-(n-k)<s\leq\min\{k,a\}$. Write $a=m+\beta,s=l+\gamma$ as in Definition 14. Then
@@
We also show two easy results about exceptional set estimate.
Lemma 25.
Recall Definition 14. Suppose we are in Type 1: $s>\min\{k,a\}$. Then there exists $A\subset\mathbb{F}_{p}^{n}$ with #$\#A\sim p^{a}$, such that
#$\#E_{s}(A;n,k)\gtrsim p^{k(n-k)}.$
@@
Lemma 26.
Recall Definition 14. Suppose we are in Type 4: $s\leq a-(n-k)$. Then for any $A\subset\mathbb{F}_{p}^{n}$ with #$\#A\geq p^{a}$, we have
#$\#E_{s}(A;n,k)=0.$
@@
3. Furstenberg set problem: upper bound
In this section, we prove Theorem 6. In other words, we will construct a $(s,t;n,k)$-set $E=\bigcup_{V\in\mathcal{V}}Y(V)$, such that
#$\#E\lesssim_{s,t,n,k}p^{\mathbf{F}(s,t;n,k)}.$
3.1. Two special cases
To get familiar with $\mathbf{F}(s,t;n,k)$, we discuss two special cases: $s=k$; $k=1$.
When $s=k$, we can morally view $Y(V)$ as $V$. Therefore, a $(k,t;n,k)$-set is of form
@@ 3.1. Two special cases
Estimating the size of $E$ is known as the “union of $k$-planes problem”. It has been resolved in finite fields by R. Oberlin [11]. The Euclidean version was resolved by the author in [3].
@@
with size as small as possible, we hope the lines in $\mathcal{L}$ are compressed in low dimensional space. Write $t=2m+\beta$, where $20\leq m\leq n-$ 2 is an integer and $\beta\in(0,2]$. Since $t\leq 2(m+1)$, we can make $\mathcal{L}$ contained in $\mathbb{F}_{p}^{m+2}$. Therefore, our goal becomes to construct a small $(s,t;m+2,1)$-set. The trick is to take Cartesian product of a $(s,\beta;2,1)$-set with $\mathbb{F}_{p}^{m}$. Let $E^{\prime}\subset\mathbb{F}_{p}^{2}$ be a $(s,\beta;2,1)$-set and $E^{\prime}$ has the form
@@
We will construct a $(s,t;m+2,1)$-set $E$ based on $E^{\prime}$. Let $\pi_{\mathbb{F}_{p}^{2}}:\mathbb{F}_{p}^{m+2}\rightarrow\mathbb{F}_{p}^{2}$ be the projection onto the first two coordinates. Define $\mathcal{L}$ to be the set of lines $L$ in $\mathbb{F}_{p}^{m+2}$ such that $\pi_{\mathbb{F}_{p}^{2}}(L)$ is a line in $\mathcal{L}^{\prime}$. Then, ##$\#\mathcal{L}\sim\#G(1,\mathbb{F}_{p}^{m+1})\cdot\#\mathcal{L}^{\prime}\sim p^{2m+\beta}=p^{t}$. If $L\in\mathcal{L}$, then $L^{\prime}=\pi_{\mathbb{F}_{p}^{2}}(L)\in\mathcal{L}^{\prime}$. We define $Y(L):=L\cap\pi_{\mathbb{F}_{p}^{2}}^{-1}(Y(L^{\prime}))$. We see that #$\#Y(L)=\#Y(L^{\prime})\sim p^{s}$, so
@@
#$\#E\leq\#E^{\prime}\cdot p^{m}\lesssim p^{m+\min\{s+\beta,\frac{3}{2}s+\frac{1}{2}\beta,s+1\}}$
By checking Definition 11, one exactly sees that
@@ 3.2. Proof of Theorem 6
If $t\geq k(n-k)$, we choose $p^{t-k(n-k)}$ points $X$ in $\mathbb{F}_{p}^{n-k}$. For each such point $x\in X$, we choose $\sim p^{k(n-k)}$ $k$-planes that pass through $x$, and require the $k$-planes to be transverse to $\mathbb{F}_{p}^{n-k}$. We obtain $\sim p^{t}$ $k$-planes which is $\mathcal{V}$. If $V$ pass through $x\in X$, we set $Y(V)=\{x\}$. We see that #$\#E=\#X\leq p^{t-k(n-k)}$.
@@
#$\#\{V\in A(k,\mathbb{F}_{p}^{n}):V\supset L\}=\#G(k-d-1,\mathbb{F}_{p}^{n-d-1})\sim p^{(k-d-1)(n-k)}.$
Let $\mathcal{V}$ be these $k$-planes. Fix $Y\subset\mathbb{F}_{p}^{d+1}$ with #$\#Y=p^{s}$. For each $V\in\mathcal{V}$, let $Y(V)=Y$. We see that #$\#E=\#Y=p^{s}$.
@@
(24) $\begin{cases}s=d+\sigma\\ t=(k-d-1)(n-k)+(d+2)m+\tau,\end{cases}$
where $d\in\{0,\dots,k-1\}$, $\sigma\in(0,1]$, $m\in\{0,\dots,n-k-1\}$, $\tau\in[0,2]$. We remark that here we allow $\tau=0$, though we assumed $\tau>$ in the canonical expression. Allowing $\tau=0$ will produce the example in Case (d).
We write $\mathbb{F}_{p}^{n}=\mathbb{F}_{p}^{n-k+d+1}\times\mathbb{F}_{p}^{k-d-1}$. Note that $\gtrsim 1$ fraction of $k$-planes intersect $\mathbb{F}_{p}^{n-k+d+1}$ at a $(d+1)$-plane. Heuristically speaking, to make #$\#E$ as small as possible, we would like $E$ to be a subset of $\mathbb{F}_{p}^{n-k+d+1}\times\mathbf{0}^{k-d-1}$. For simplicity, we write $\mathbb{F}_{p}^{n-k+d+1}\times\mathbf{0}^{k-d-1}$ as $\mathbb{F}_{p}^{n-k+d+1}$, thinking of it as a subspace of $\mathbb{F}_{p}^{n}$. This will not cause any ambiguity.
@@
[…]
is #$\sim\#G(k-d-1,\mathbb{F}_{p}^{n-k})\sim p^{(k-d-1)(n-k)}$. We just need to find $\mathcal{W}\subset A(d+1,\mathbb{F}_{p}^{n-k+d+1})$ with #$\#\mathcal{W}\sim p^{t-(k-d-1)(n-k)}=p^{(d+2)m+\tau}$, and for each $W\in\mathcal{W}$, to find $Y(W)\subset W$ with #$\#Y(W)\sim p^{s}$, such that
(25) #$\#\bigcup_{W\in\mathcal{W}}Y(W)\lesssim p^{\mathbf{F}(s,t;n,k)}.$
If this is done, then we let $\mathcal{V}$ be those $k$-planes $V$ that contain some $W\in\mathcal{W}$, and $Y(V)=Y(W)$ if $V\supset W$. We see that #$\#\mathcal{V}\sim\#\{W\}\cdot p^{(k-d-1)(n-k)}\sim p^{t}$, and hence $E=\bigcup_{V\in\mathcal{V}}Y(V)$ is a $(s,t;n,k)$-set. Also, we have the upper bound
#$\#E=\#\bigcup_{W\in\mathcal{W}}Y(W)\lesssim p^{\mathbf{F}(s,t;n,k)}.$
It suffices to find a $(s,(d+2)m+\tau;n-k+d+1,d+1)$-set $E_{1}=\bigcup_{W\in\mathcal{W}}Y(W)$ such that (25) holds. Note that #$\#A(d+1,\mathbb{F}_{p}^{m+d+2})\sim p^{(d+2)(m+1)}$ and $(d+2)m+\tau\leq(d+2)(m+1)$. We will construct the $(d+1)$-planes $\mathcal{W}$ so that they lie in $\mathbb{F}_{p}^{m+d+2}$. It suffices to find a $(s,(d+2)m+\tau;m+d+2,d+1)$-set $E_{1}=\bigcup_{W\in\mathcal{W}}Y(W)$ such that (25) holds.
@@
#$\#E_{2}\lesssim p^{\min\{\sigma+\tau,\frac{3}{2}\sigma+\frac{1}{2}\tau,\sigma+1\}}.$
The existence of $E_{2}$ was shown in Proposition 13.
Next, we define $\mathcal{U}:=\{L\times\mathbb{F}_{p}^{d}:L\in\mathcal{L}\}$ which is a subset of $A(d+1,\mathbb{F}_{p}^{d+2})$. And for each $U=L\times\mathbb{F}_{p}^{d}\in\mathcal{U}$, define $Y(U)=Y(L)\times\mathbb{F}_{p}^{d}$. We see that #$\#\mathcal{U}=\#\mathcal{L}\sim p^{\tau}$, #$\#Y(U)=\#Y(L)\cdot p^{d}\sim p^{\sigma+d}=p^{s}$. We also see that $E_{3}:=\bigcup_{U\in\mathcal{U}}Y(U)$ is contained in $E_{2}\times\mathbb{F}_{p}^{d}$, and hence
#$\#E_{3}\lesssim p^{d+\min\{\sigma+\tau,\frac{3}{2}\sigma+\frac{1}{2}\tau,\sigma+1\}}.$
Finally, we define $\mathcal{W}$. For each $U\in\mathcal{U}$, consider the $(m+d+1)$-plane $U\times\mathbb{F}_{p}^{m}$. Let $\pi_{U}:U\times\mathbb{F}_{p}^{m}\rightarrow U$ be the projection (naturally defined). We let $\mathcal{W}_{U}$ be the set of $(d+1)$-planes $W$ contained in $U\times\mathbb{F}_{p}^{m}$ such that $\pi_{U}(W)=U$. Note that $\gtrsim 1$ fraction of the $(d+1)$-planes in $U\times\mathbb{F}_{p}^{m}$ lie in $\mathcal{W}_{U}$, so #$\#\mathcal{W}_{U}\sim\#A(d+1,\mathbb{F}_{p}^{m+d+1})\sim p^{(d+2)m}$.
We define $\mathcal{W}=\bigcup_{U\in\mathcal{U}}\mathcal{W}_{U}$ (actually this is a disjoint union). For $W\in\mathcal{W}_{U}$, define $Y(W)=W\cap\pi_{U}^{-1}(Y(U))$. We see that #$\#\mathcal{W}\sim\#\mathcal{U}\cdot p^{(d+2)m}\sim p^{(d+2)m+\tau}$, and #$\#Y(W)=\#Y(U)\sim p^{s}$. Hence, $E_{1}=\bigcup_{W\in\mathcal{W}}Y(W)$ is a $(s,(d+2)m+\tau;m+d+2,d+1)$-set. Since $E_{1}\subset E_{3}\times\mathbb{F}_{p}^{m}$, we have
#$\#E_{1}\leq\#E_{3}\cdot p^{m}\lesssim p^{d+m+\min\{\sigma+\tau,\frac{3}{2}\sigma+\frac{1}{2}\tau,\sigma+1\}}.$
@@
We still use (24), but the range of $\tau$ is $\tau\in(2,d+2]$. Note the worst case is when $\tau=d+2$. What we need to do is, for
(26) $\begin{cases}s=d+\sigma\\ t=(k-d-1)(n-k)+(d+2)(m+1),\end{cases}$
@@
#$\#E\lesssim p^{s+m+1}.$
The trick is to still use the construction in Case (c), with $\tau$ replaced by 00 and $m$ replaced by $m+1$. Therefore, we can construct a $(s,t;n,k)$-set $E$ such that
#$\#E\lesssim p^{d+m+1+\min\{\sigma+0,\frac{3}{2}\sigma+\frac{1}{2}0,\sigma+1\}=p^{d+m+1+\sigma}=p^{s+m+1}}.$
4. Furstenberg set problem: Lower bound I
To get more familiar with the problem, we first prove two special cases of Theorem 7: $s=0$; $t=0$.
Proposition 27.

### v2

@@ 2.2 Useful lemmas
To show this, we just need to prove that for a fixed $a\in(0,n]$, the ranges of $s$ determined by four cases form a partition of $(0,\infty)$. Type 1 corresponds to $s\in(\min\{a,k\},\infty)$. Type 4 corresponds to $s\in(0,a-(n-k)]$. Note that $a-(n-k)\leq\min\{a,k\}$, so we just need to show the ranges of $s$ determined by Type 2 and Type 3 from a partition of $(a-(n-k),\min\{a,k\}]$. Since the ranges of $s$ determined by Type2 and Type 3 are disjoint, we just need to show
$\{s:s\textup{~satisfies~Type~}2\}\bigcup\{s:s\textup{~satisfies~Type~}3\}=(a-(n-k),\min\{a,k\}].$
We first show $\subset$". Suppose $s$ satisfies the condition in Type 2. We want to show s>a(nk)s>a-(n-k), which is equivalent to $m<n+l-k+\gamma-\beta$. This is true since $m\leq n+l-k$ and $\gamma>\beta$. Suppose $s$ satisfies the condition in Type 3. Again, we want to show $m<n+l-k+\gamma-\beta$. This is true since $m\leq n+l-k-1$ and $\gamma-\beta>-\beta\geq-1$.
Next, we show $\supset$”. Suppose $s=l+\gamma\in(a-(n-k),\min\{a,k\}]$. s>a(nk)s>a-(n-k) implies
@@
Lemma 2.14.
$\mathbf{M}(a,s;n,k)$ is decreasing in $a$ and increasing in $s$.
Lemma 2.15.
@@
Let $a=m+\beta,s=l+\gamma$ be the canonical expression. We just check four types as in Definition 2.4.
@@
Note that the integer part $(m,l)$ is reduced to (n1,l)(n-1,l), so we can continue doing the iteration.
@@
Lemma 2.16.
Suppose $a-(n-k)<s\leq\min\{k,a\}$. Write $a=m+\beta,s=l+\gamma$ as in Definition 2.4. Then
@@
We also show two easy results for the exceptional set estimate.
Lemma 2.17.
Recall Definition 2.4. Suppose that we are in Type 1: $s>\min\{k,a\}$. Then there exists $A\subset\mathbb{F}_{p}^{n}$ with $A\sim p^{a}$, such that
$E_{s}(A;n,k)\gtrsim p^{k(n-k)}.$
@@
Lemma 2.18.
Recall Definition 2.4. Suppose that we are in Type 4: $s\leq a-(n-k)$. Then for any $A\subset\mathbb{F}_{p}^{n}$ with $A\geq p^{a}$, we have
$E_{s}(A;n,k)=0.$
@@
3 Furstenberg set problem: upper bound
In this section, we prove Theorem 1.7. In other words, we will construct a $(s,t;n,k)$-set $E=\bigcup_{V\in\mathcal{V}}Y(V)$, such that
$E\lesssim_{s,t,n,k}p^{\mathbf{F}(s,t;n,k)}.$
3.1 Two special cases
To familiarize ourselves with $\mathbf{F}(s,t;n,k)$, we discuss two special cases: $s=k$; $k=1$.
When $s=k$, we can morally view $Y(V)$ as $V$. Therefore, a $(k,t;n,k)$-set is of the form
@@ 3.1 Two special cases
Estimating the size of $E$ is known as the problem of “union of $k$ planes”. It has been resolved in finite fields by R. Oberlin [15]. The Euclidean version was resolved by the author in [7].
@@
with size as small as possible, we hope that the lines in $\mathcal{L}$ are compressed in a low-dimensional space. Write $t=2m+\beta$, where $0\leq m\leq n-2$ is an integer, and $\beta\in(0,2]$. Since $t\leq 2(m+1)$, we can make $\mathcal{L}$ contained in $\mathbb{F}_{p}^{m+2}$. Therefore, our goal becomes to construct a small $(s,t;m+2,1)$-set. The trick is to take the Cartesian product of a $(s,\beta;2,1)$-set with $\mathbb{F}_{p}^{m}$. Let $E^{\prime}\subset\mathbb{F}_{p}^{2}$ be a $(s,\beta;2,1)$-set and let $E^{\prime}$ have the form
@@
We will construct a $(s,t;m+2,1)$-set $E$ based on $E^{\prime}$. Let $\pi_{\mathbb{F}_{p}^{2}}:\mathbb{F}_{p}^{m+2}\rightarrow\mathbb{F}_{p}^{2}$ be the projection onto the first two coordinates. Define $\mathcal{L}$ to be the set of lines $L$ in $\mathbb{F}_{p}^{m+2}$ such that $\pi_{\mathbb{F}_{p}^{2}}(L)$ is a line in $\mathcal{L}^{\prime}$. Then, $\mathcal{L}\sim\#G(1,\mathbb{F}_{p}^{m+1})\cdot\#\mathcal{L}^{\prime}\sim p^{2m+\beta}=p^{t}$. If $L\in\mathcal{L}$, then $L^{\prime}=\pi_{\mathbb{F}_{p}^{2}}(L)\in\mathcal{L}^{\prime}$. We define $Y(L):=L\cap\pi_{\mathbb{F}_{p}^{2}}^{-1}(Y(L^{\prime}))$. We see that $Y(L)=\#Y(L^{\prime})\sim p^{s}$, so
@@
$E\leq\#E^{\prime}\cdot p^{m}\lesssim p^{m+\min\{s+\beta,\frac{3}{2}s+\frac{1}{2}\beta,s+1\}}$
By checking Definition 2.1, one exactly sees that
@@
[…]
If $t\geq k(n-k)$, we choose $p^{t-k(n-k)}$ points $X$ in $\mathbb{F}_{p}^{n-k}$. For each such point $x\in X$, we choose $\sim p^{k(n-k)}$ $k$-planes that pass through $x$, and require the $k$-planes to be transverse to $\mathbb{F}_{p}^{n-k}$. We obtain $\sim p^{t}$ $k$-planes which is $\mathcal{V}$. If $V$ passes through $x\in X$, we set $Y(V)=\{x\}$. We see that $E=\#X\leq p^{t-k(n-k)}$.
@@
$\{V\in A(k,\mathbb{F}_{p}^{n}):V\supset L\}=\#G(k-d-1,\mathbb{F}_{p}^{n-d-1})\sim p^{(k-d-1)(n-k)}.$
Let $\mathcal{V}$ be these $k$-planes. Fix $Y\subset\mathbb{F}_{p}^{d+1}$ with $Y=p^{s}$. For each $V\in\mathcal{V}$, let $Y(V)=Y$. We see that $E=\#Y=p^{s}$.
@@
$\begin{cases}s=d+\sigma\\ t=(k-d-1)(n-k)+(d+2)m+\tau,\end{cases}$ (24)
where $d\in\{0,\dots,k-1\}$, $\sigma\in(0,1]$, $m\in\{0,\dots,n-k-1\}$, $\tau\in[0,2]$. We remark that here we allow $\tau=0$, although we assumed $\tau>$ in the canonical expression. Allowing $\tau=0$ will produce the example in Case (d).
We write $\mathbb{F}_{p}^{n}=\mathbb{F}_{p}^{n-k+d+1}\times\mathbb{F}_{p}^{k-d-1}$. Note that $\gtrsim 1$ fraction of $k$-planes intersect $\mathbb{F}_{p}^{n-k+d+1}$ at a $(d+1)$-plane. Heuristically speaking, to make #E\#E as small as possible, we would like $E$ to be a subset of $\mathbb{F}_{p}^{n-k+d+1}\times\mathbf{0}^{k-d-1}$. For simplicity, we write $\mathbb{F}_{p}^{n-k+d+1}\times\mathbf{0}^{k-d-1}$ as $\mathbb{F}_{p}^{n-k+d+1}$, thinking of it as a subspace of $\mathbb{F}_{p}^{n}$. This will not cause any ambiguity.
@@
$\{V\in A(k,\mathbb{F}_{p}^{n}):V\supset W,V\textup{~is~transverse~to~}\mathbb{F}_{p}^{n-k+d+1}\}$
is $\sim\#G(k-d-1,\mathbb{F}_{p}^{n-k})\sim p^{(k-d-1)(n-k)}$. We just need to find $\mathcal{W}\subset A(d+1,\mathbb{F}_{p}^{n-k+d+1})$ with $\mathcal{W}\sim p^{t-(k-d-1)(n-k)}=p^{(d+2)m+\tau}$, and for each $W\in\mathcal{W}$, to find $Y(W)\subset W$ with $Y(W)\sim p^{s}$, such that
$\bigcup_{W\in\mathcal{W}}Y(W)\lesssim p^{\mathbf{F}(s,t;n,k)}.$ (25)
If this is done, then we let $\mathcal{V}$ be those $k$-planes $V$ that contain some $W\in\mathcal{W}$, and $Y(V)=Y(W)$ if $V\supset W$. We see that $\mathcal{V}\sim\#\{W\}\cdot p^{(k-d-1)(n-k)}\sim p^{t}$, and hence $E=\bigcup_{V\in\mathcal{V}}Y(V)$ is a $(s,t;n,k)$-set. Also, we have the upper bound
$E=\#\bigcup_{W\in\mathcal{W}}Y(W)\lesssim p^{\mathbf{F}(s,t;n,k)}.$
It suffices to find a $(s,(d+2)m+\tau;n-k+d+1,d+1)$-set $E_{1}=\bigcup_{W\in\mathcal{W}}Y(W)$ such that (25) holds. Note that $A(d+1,\mathbb{F}_{p}^{m+d+2})\sim p^{(d+2)(m+1)}$ and $(d+2)m+\tau\leq(d+2)(m+1)$. We will construct the $(d+1)$-planes $\mathcal{W}$ so that they lie in $\mathbb{F}_{p}^{m+d+2}$. It suffices to find a $(s,(d+2)m+\tau;m+d+2,d+1)$-set $E_{1}=\bigcup_{W\in\mathcal{W}}Y(W)$ such that (25) holds.
@@
$E_{2}\lesssim p^{\min\{\sigma+\tau,\frac{3}{2}\sigma+\frac{1}{2}\tau,\sigma+1\}}.$
The existence of $E_{2}$ was shown in Proposition 2.3.
Next, we define $\mathcal{U}:=\{L\times\mathbb{F}_{p}^{d}:L\in\mathcal{L}\}$ which is a subset of $A(d+1,\mathbb{F}_{p}^{d+2})$. And for each $U=L\times\mathbb{F}_{p}^{d}\in\mathcal{U}$, define $Y(U)=Y(L)\times\mathbb{F}_{p}^{d}$. We see that $\mathcal{U}=\#\mathcal{L}\sim p^{\tau}$, $Y(U)=\#Y(L)\cdot p^{d}\sim p^{\sigma+d}=p^{s}$. We also see that $E_{3}:=\bigcup_{U\in\mathcal{U}}Y(U)$ is contained in $E_{2}\times\mathbb{F}_{p}^{d}$, and hence
$E_{3}\lesssim p^{d+\min\{\sigma+\tau,\frac{3}{2}\sigma+\frac{1}{2}\tau,\sigma+1\}}.$
Finally, we define $\mathcal{W}$. For each $U\in\mathcal{U}$, consider the $(m+d+1)$-plane $U\times\mathbb{F}_{p}^{m}$. Let $\pi_{U}:U\times\mathbb{F}_{p}^{m}\rightarrow U$ be the projection (naturally defined). We let $\mathcal{W}_{U}$ be the set of $(d+1)$-planes $W$ contained in $U\times\mathbb{F}_{p}^{m}$ such that $\pi_{U}(W)=U$. Note that $\gtrsim 1$ fraction of the $(d+1)$-planes in $U\times\mathbb{F}_{p}^{m}$ lie in $\mathcal{W}_{U}$, so $\mathcal{W}_{U}\sim\#A(d+1,\mathbb{F}_{p}^{m+d+1})\sim p^{(d+2)m}$.
We define $\mathcal{W}=\bigcup_{U\in\mathcal{U}}\mathcal{W}_{U}$ (actually this is a disjoint union). For $W\in\mathcal{W}_{U}$, define $Y(W)=W\cap\pi_{U}^{-1}(Y(U))$. We see that $\mathcal{W}\sim\#\mathcal{U}\cdot p^{(d+2)m}\sim p^{(d+2)m+\tau}$, and $Y(W)=\#Y(U)\sim p^{s}$. Hence, $E_{1}=\bigcup_{W\in\mathcal{W}}Y(W)$ is a $(s,(d+2)m+\tau;m+d+2,d+1)$-set. Since $E_{1}\subset E_{3}\times\mathbb{F}_{p}^{m}$, we have
$E_{1}\leq\#E_{3}\cdot p^{m}\lesssim p^{d+m+\min\{\sigma+\tau,\frac{3}{2}\sigma+\frac{1}{2}\tau,\sigma+1\}}.$
@@
We still use (24), but the range of $\tau$ is $\tau\in(2,d+2]$. Note that the worst case is when $\tau=d+2$. What we need to do is for
$\begin{cases}s=d+\sigma\\ t=(k-d-1)(n-k)+(d+2)(m+1),\end{cases}$ (26)
@@
$E\lesssim p^{s+m+1}.$
The trick is to still use the construction in Case (c), with $\tau$ replaced by 0 and $m$ replaced by $m+1$. Therefore, we can construct a $(s,t;n,k)$-set $E$ such that
$E\lesssim p^{d+m+1+\min\{\sigma+0,\frac{3}{2}\sigma+\frac{1}{2}0,\sigma+1\}=p^{d+m+1+\sigma}=p^{s+m+1}}.$
4 Furstenberg set problem: Lower bound I
To become more familiar with the problem, we first prove two special cases of Theorem 1.8: $s=0$; $t=0$.
Proposition 4.1.